### v3

Abstract
Homomorphic ring signature schemes combine the strong anonymity of ring signatures with the computability of homomorphic signatures, demonstrating significant potential in scenarios requiring both anonymous data provenance and verifiable homomorphic computation (e.g., confidential blockchain transactions and secure multi-party computation). However, no feasible homomorphic ring signature scheme currently exists.
In this work, we propose the first lattice-based linearly homomorphic ring signature scheme. Proven secure in the standard model under the small integer solution (SIS) assumption, our scheme achieves strong anonymity under full key exposure and unforgeability against insider corruption attacks. As the first unified framework for ring signatures and linear homomorphic signatures, this construction provides a post-quantum-secure solution for the aforementioned applications, advancing the development of privacy-enhanced homomorphic computation.
keywords:
Lattice, homomorphic ring signature, anonymity, unforgeability, Short Integer Solution (SIS)
1 Introduction
Functioning as a fundamental pillar of public-key cryptography, the digital signature mechanism was formally presented by Diffie and Hellman in their landmark 1976 paper, New Directions in Cryptography [1]. This seminal work catalyzed extensive research into digital signatures. Driven by objectives to enhance efficiency and fortify security, subsequent research has engendered a sophisticated body of theory and an array of practical techniques. Representative works can be found in references [2, 3, 4, 5, 6].
In 2001, Rivest, Shamir, and Tauman [7, 8] first proposed the ring signature scheme. The core concept of this cryptographic primitive is as follows: the signer can select a group of users through their public keys and then sign a message on behalf of this group (referred to as a ”ring”). Signing in the name of these users implies that it is impossible to determine which specific user within the ring actually signed the message. Ring signatures exhibit notable flexibility characteristics: the signer can freely and spontaneously form the ring structure without the involvement of a trusted third party; in fact, ring members do not even need to be aware of each other’s existence. Ring signature technology has garnered sustained attention in academia due to its potential application value in addressing real-world technical and social challenges. For instance, it protects whistleblower identities [7], enables participant anonymity in electronic voting systems and digital cash protocols [9], facilitates transaction anonymization in blockchain applications [10], and particularly in the Monero cryptocurrency system, effectively achieves transaction untraceability [11]. Representative ring signature schemes can be referred to in [12, 13, 14, 15, 16].
Homomorphic signatures were initially formalized by Johnson et al. [17] in 2002 as a cryptographic primitive that permits authorized computations on authenticated data while maintaining verifiability without access to the original signing key. This functionality has demonstrated significant utility across various applications, most notably in network coding protocols [18, 19, 20] and cloud-based outsourced computation systems [21, 22, 23]. Subsequent research has systematically organized homomorphic signature schemes into three principal classes: linearly homomorphic schemes supporting additive operations [24, 25, 26, 27, 28], polynomially homomorphic schemes enabling polynomial function evaluation [29, 30, 31, 32], and fully homomorphic schemes allowing arbitrary computations [33, 34, 35].
Although ring signatures and homomorphic signatures each offer robust cryptographic capabilities—namely, anonymity and the ability to compute on authenticated data, respectively—their combination remains largely unexplored in the literature. A homomorphic ring signature scheme would inherently support both the anonymity guarantees of ring signatures and the malleability of homomorphic signatures, enabling computation on signed data while preserving signer privacy within dynamically formed rings. Such a scheme could unlock novel applications in privacy-preserving distributed systems, such as confidential blockchain transactions requiring signed input aggregation or secure multi-party computation scenarios where participants demand both anonymity and verifiable correctness. However, designing a practical homomorphic ring signature scheme presents significant challenges, particularly in balancing security, efficiency, and functionality.
Our contributions. In this study, we address this gap by proposing the first lattice-based linearly homomorphic ring signature (LHRS) scheme , achieving quantum-resistant security while preserving the anonymity and computational practicality of the two underlying primitives.
Organization. The structure of this paper is as follows: Section 2 introduces the basic notation definitions and theoretical foundations. Section 3 presents the definition of the signature scheme and its security model. Section 4 elaborates on the construction of the scheme and provides a formal correctness proof. Section 5 demonstrates that the proposed scheme satisfies unconditional anonymity and existential unforgeability. Finally, Section 6 concludes the paper and discusses open problems.
2 Preliminaries
2.1 Notation
Mathematical Notation: Vectors are denoted by bold lowercase letters (e.g., $\mathbf{u}$, $\mathbf{v}$), and matrices are denoted by uppercase letters (e.g., $A$, $T$). The horizontal concatenation of matrices is denoted as $A\mid T]$ or $A\parallel T]$. The ring of integers is denoted by $\mathbb{Z}$, the ring of integers modulo $q$ by $\mathbb{Z}_{q}$ (or $\mathbb{Z}/q\mathbb{Z}$), a general field by $\mathbb{F}$, and the finite field with $q$ elements (where $q$ is a prime or a prime power) by $\mathbb{F}_{q}$. The $n$-dimensional Euclidean space is denoted by $\mathbb{R}^{n}$.
Matrix Norms and Orthogonalization: For a matrix $A=[\mathbf{a}_{1},\dots,\mathbf{a}_{m}]\in\mathbb{R}^{n\times m}$, the matrix norm is defined as $\|A\|=\max_{1\leq i\leq m}\|\mathbf{a}_{i}\|$, where $\|\mathbf{a}_{i}\|$ represents the $\ell_{2}$-norm of the vector $\mathbf{a}_{i}$. Furthermore, let $\widetilde{A}=[\widetilde{\mathbf{a}}_{1},\dots,\widetilde{\mathbf{a}}_{n}]$ denote the Gram-Schmidt orthogonalization of $A$, i.e.,
$\widetilde{\mathbf{a}}_{1}=\mathbf{a}_{1},\quad\widetilde{\mathbf{a}}_{i}=\mathbf{a}_{i}-\sum_{j=1}^{i-1}\frac{\langle\mathbf{a}_{i},\widetilde{\mathbf{a}}_{j}\rangle}{\langle\widetilde{\mathbf{a}}_{j},\widetilde{\mathbf{a}}_{j}\rangle}\widetilde{\mathbf{a}}_{j},\quad 2\leq i\leq n,$
where $\langle\cdot,\cdot\rangle$ denotes the standard inner product in Euclidean space.
Asymptotic Notation: For asymptotic analysis, we employ the following notation:
  • $f(n)=\mathcal{O}(g(n))$ if $\exists c>0,N\in\mathbb{N}$ such that $\forall n>N$, $f(n)\leq cg(n)$
  • $f(n)=\widetilde{\mathcal{O}}(g(n))$ when $f(n)=\mathcal{O}(g(n)\log^{c^{\prime}}n)$ for some $c^{\prime}>0$
  • $f(n)=\omega(g(n))$ if $\forall c>0$, $\exists N$ such that $g(n)\leq cf(n)$ whenever $n>N$
Let $n$ be the security parameter. A function $f(n)$ is said to be polynomial (written $f(n)=\mathrm{poly}(n)$) if there exists $c>0$ such that $f(n)=\mathcal{O}(n^{c})$; conversely, it is negligible (denoted $\mathrm{negl}(n)$) if for every $c>0$ we have $f(n)=\mathcal{O}(n^{-c})$. An event is said to occur with overwhelming probability when its probability satisfies $\Pr[\text{event}]\geq 1-\mathrm{negl}(n)$.
Probabilistic Notation and Algorithms: For any given distribution $\mathcal{D}$, the expression $x\sim\mathcal{D}$ signifies that the variable $x$ follows the distribution $\mathcal{D}$, while $x\leftarrow\mathcal{D}$ denotes the process of sampling a random value from $\mathcal{D}$. When considering a finite set $\mathcal{X}$, the notation $x\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathcal{X}$ indicates that $x$ is chosen uniformly at random from $\mathcal{X}$. Regarding probabilistic polynomial-time (PPT) algorithms $\mathsf{Alg}$, the notation $y\leftarrow\mathsf{Alg}(x)$ represents the procedure of generating an output $y$ by executing the algorithm on input $x$.
2.2 Statistical Distance, Entropy and Decomposition Algorithm
Definition 1.
(Statistical distance, [36]) Let $M\subset\mathbb{R}^{n}$ be a finite or countable set, and let $X$ and $Y$ be discrete random variables taking values in $M$. The statistical distance $X$ and $Y$ is defined as:
$\Delta(X,Y)=\frac{1}{2}\sum_{a\in M}\left|\mathbb{P}\{X=a\}-\mathbb{P}\{Y=a\}\right|.$
Two probability distributions $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$ are said to be statistically indistinguishable if their statistical distance
$\Delta(\mathcal{D}_{0},\mathcal{D}_{1})$
is a negligible function in the security parameter $n$.
Definition 2.
( Min-entropy[33]) For a random variable $X$, its min-entropy is defined as:
$H_{\infty}(X)=-\log\left(\max_{x\in X}\Pr\{X=x\}\right).$
The average conditional min-entropy of a random variable $X$ conditional on a correlated variable $Y$ is defined as:
$H_{\infty}(X|Y)=-\log\left(\mathbb{E}_{y\in Y}\left\{\max_{x\in X}\Pr\{X=x|Y=y\}\right\}\right).$
The optimal probability of an unbounded adversary guessing $X$ given the correlated value $Y$ is $2^{-H_{\infty}(X|Y)}$[33].
Lemma 1.
([33]) Let $X$, $Y$ be arbitrarily random variables where the support of $Y$ lies in $\mathcal{Y}$, Then
$H_{\infty}(X|Y)>H_{\infty}(X)-\log(|\mathcal{Y}|).$
The following decomposition algorithm is derived from [24]. This algorithm decomposes a vector $\mathbf{w}\in\mathbb{F}_{2}^{2k}$ into a pair of vectors $(\mathbf{u},\mathbf{v})$ belonging to a specific set, as described below:
Theorem 1.
( Decomposition Algorithm[24]) Let $k$ be an odd integer with $k=\mathrm{poly}(n)$. Define the sets
$\mathcal{X}$ $=\big\{x\in\mathbb{F}_{2}^{2k}\mid\|x\|=k-1\big\}\sqcup\{0\},$
$\mathcal{Y}$ $=\big\{y\in\mathbb{F}_{2}^{2k}\mid\|y\|=k\big\}\sqcup\{0\},$
$\mathcal{Z}$ $=\big\{z\in\mathbb{F}_{2}^{2k}\mid\|z\|=k+1\big\}\sqcup\{0\}.$
There exists a polynomial-time algorithm $\textsf{Decompose}(\mathbf{w})$ that, given an input $\mathbf{w}\in\mathbb{F}_{2}^{2k}$, outputs a pair of vectors $(\mathbf{u},\mathbf{v})$ such that:
  • $\mathbf{u}\in\mathcal{X}$ and $\mathbf{v}\in\mathcal{Y}$ or $\mathcal{Z}$;
  • if $\mathbf{w}\neq 0$, then at most one of $\mathbf{u}$ and $\mathbf{v}$ is zero.
2.3 Background on Lattices and Hard Problems
Definition 3.
(Lattice [37]) Let $\Lambda\subset\mathbb{R}^{n}$ be a non-empty subset. $\Lambda$ is called a lattice if:
(1) it is an additive subgroup of $\mathbb{R}^{n}$;
(2) there exists a positive constant $\lambda=\lambda(\Lambda)>0$ such that
$\mathrm{min}\{\parallel\mathbf{x}\parallel|\mathbf{x}\in\Lambda,\mathbf{x}\neq 0\}=\lambda.$
$\lambda$ is called the minimum distance.
A full-rank lattice $\Lambda\subseteq\mathbb{R}^{n}$ can be equivalently characterized by a basis $B=\{\mathbf{b}_{1},\dots,\mathbf{b}_{n}\}$, where each $\mathbf{b}_{i}\in\mathbb{R}^{n}$ is linearly independent. The lattice generated by $B$ comprises all integer linear combinations of the basis vectors:
$\Lambda(B)=\left\{\sum_{i=1}^{n}c_{i}\mathbf{b}_{i}\;\Big|\;\bm{c}=(c_{1},\dots,c_{n})^{\top}\in\mathbb{Z}^{n}\right\}.$
The dual lattice $\Lambda^{*}$, which is intrinsically associated with $\Lambda$, is defined as the set of vectors satisfying certain orthogonality conditions with respect to $\Lambda$:
$\Lambda^{*}=\left\{\mathbf{v}\in\mathbb{R}^{n}\;\Big|\;\langle\mathbf{v},\mathbf{u}\rangle\in\mathbb{Z}\text{ for all }\mathbf{u}\in\Lambda\right\}.$
Definition 4.
($q$-ary lattices[37]) Let $A\in\mathbb{Z}_{q}^{n\times m}$, $\mathbf{u}\in\mathbb{Z}^{n}$. The following two $q$-ary lattices are defined as:
[…]
The set $\Lambda_{q}^{\mathbf{u}}$ is a coset of $\Lambda_{q}^{\bot}$ since $\Lambda_{q}^{\mathbf{u}}=\Lambda_{q}^{\bot}+\mathbf{x}$ for any $\mathbf{x}$ such that $A\cdot\mathbf{x}\equiv\mathbf{u}(\mathrm{mod}q)$.
Definition 5.
(Short integer solution[37]) Let $n$, $m$, $q$ be positive integers, with $m=\mathrm{poly(}n)$. Let $A\in\mathbb{Z}_{q}^{n\times m}$ be a uniformly distributed random matrix over $\mathbb{Z}_{q}$, and let $\beta\in\mathbb{R}$ such that $0<\beta<q$. The $\mathrm{SIS}$ problem is to find a short integer solution $\mathbf{x}$ satisfying the following condition:
$A\cdot\mathbf{x}\equiv 0(\mathrm{mod}q),\quad\mathrm{and}\quad\mathbf{x}\neq 0,\parallel\mathbf{x}\parallel\leq\beta.$
We write the above $\mathrm{SIS}$ problem as $\mathrm{SIS}_{q,n,m,\beta}$ or $\mathrm{SIS}_{q,\beta}$.
Theorem 2.
(Worst-case to average-case reduction [38]) For any polynomial bounded $m=\text{poly}(n)$, and any $\beta>0$, if $q\geq\beta\cdot\omega(\sqrt{n\log n})$, then solving the average-case problem $\mathrm{SIS}_{q,\beta}$ is at least as hard as solving the worst-case problem $\mathrm{SIVP}{\gamma}$ on any $n$-dimensional lattice for $\gamma=\beta\cdot\tilde{O}(\sqrt{n})$.
2.4 Gaussian Distribution and Its Related Algorithms
Definition 6.
( Discrete Gaussian distributions[36]) Let $s$ be a positive real number and $\mathbf{c}\in\mathbb{R}^{n}$ be a vector. The Gaussian function centered at $\mathbf{c}$ with parameter $s$ is defined as: $\rho_{s,\mathbf{c}}(\mathbf{x})=e^{\frac{-\pi}{s^{2}}\parallel\mathbf{x}-\mathbf{c}\parallel^{2}}.$ The discrete Gaussian measure $\mathcal{D}_{\Lambda,s,\mathbf{c}}$ defined on the lattice $\Lambda$ is given by:
$\mathcal{D}_{\Lambda,s,\mathbf{c}}=\frac{\rho_{s,\mathbf{c}}(\mathbf{x})}{\rho_{s,\mathbf{c}}(\Lambda)},$
where $\rho_{s,\mathbf{c}}(\Lambda)=\sum_{\mathbf{x}\in\Lambda}\rho_{s,\mathbf{c}}(\mathbf{x})$.
Theorem 3.
([39]) Let $q\geq 3$ be odd, $n$ be a positive integer, and let $m:=\lceil 6n\log q\rceil.$ There is a probabilistic polynomial-time algorithm TrapGen(q,n,m)(q,n,m) that outputs a pair $(A,T)$ such that $A$ is statistically close to a uniform rank $n$ matrix in $\mathbb{Z}_{q}^{n\times m}$, and $T\in\mathbb{Z}^{n\times n}$ is a basis for $\Lambda_{q}^{\bot}(A)$ satisfying
$\parallel\widetilde{T}\parallel\leq\mathcal{O}(\sqrt{n\log q})\quad\text{and}\quad\parallel T\parallel\leq\mathcal{O}(n\log q).$
In [38], Gentry and colleagues presented a method for sampling from discrete Gaussian distributions with any short basis.
Lemma 2.
(Sampling from discrete Gaussian [38]) Let $q\geq 2$ , $A\in\mathbb{Z}_{q}^{n\times m}$ with $m>n$ and let $T$ be a basis for $\Lambda_{q}^{\bot}(A)$ and $s\geq\widetilde{T}\cdot\omega(\sqrt{\log m})$. Then for $\mathbf{c}\in\mathbb{R}^{n}$ and $\mathbf{u}\in\mathbb{Z}_{q}^{n}$:
  1. 1.
[…]
  2. 2.
[…]
For any parameter $s$ such that $s\geq\omega(\sqrt{\log n})$, the probabilistic algorithm $\mathsf{SampleDom}(1^{n},s)$ produces samples following a discrete Gaussian distribution over $\mathbb{Z}^{n}$. Given a randomly sampled vector $\mathbf{x}\leftarrow\mathsf{SampleDom}(1^{n},s)$, its output distribution is statistically close to $\mathcal{D}_{\mathbb{Z}^{n},s}$. Additionally, the conditional min-entropy of the generated samples is $\omega(\log n)$ [38].
Lemma 3.
([38]) Let $n$ and $h$ be positive integers, and let $q$ be a prime satisfying $n\geq 2h\log q$. Then, for all matrices $A\in\mathbb{Z}_{q}^{h\times n}$ except a negligible fraction of $2q^{-h}$, and for any $s\geq\omega(\sqrt{\log n})$, the distribution of $\mathbf{\alpha}=A\cdot\mathbf{x}\pmod{q}$ is statistically indistinguishable from uniform over $\mathbb{Z}_{q}^{h}$, where $\mathbf{x}$ is sampled as $\mathbf{x}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\textsf{SampleDom}(1^{n},s)$.
The generalized sampling algorithm $\mathsf{GenSamplePre}$ was first proposed in [41]. [15] extended this technique by selecting different parameters and expanding the lattice structure, obtaining the following improved lattice extension technique that is more suitable for the subsequent ring signature scheme.
Theorem 4.
(Sampling Preimage for Extended Lattice[15]) Let $k,k_{1},k_{2},k_{3},k_{4}$ be positive integers with $k=k_{1}+k_{2}+k_{3}+k_{4}$. Let $S=[k]=\{1,2,\ldots,k\}$ (interpretable as an index set), and denote $A_{S}=[A_{S_{1}}|A_{S_{2}}|A_{S_{3}}|A_{S_{4}}]$, where $A_{S_{1}}\in\mathbb{Z}_{q}^{h\times k_{1}n}$, $A_{S_{2}}\in\mathbb{Z}_{q}^{h\times k_{2}n}$, $A_{S_{3}}\in\mathbb{Z}_{q}^{h\times k_{3}n}$, and $A_{S_{4}}\in\mathbb{Z}_{q}^{h\times k_{4}n}$. Let $A_{R}=[A_{S_{1}}|A_{S_{3}}]\in\mathbb{Z}_{q}^{h\times(k_{1}+k_{3})n}$, given a short basis $T_{R}$ for the lattice $\Lambda_{q}^{\perp}(A_{R})$, and a real number $V\geq\|\widetilde{T}_{R}\|\omega(\sqrt{\log(k_{1}+k_{3})n})$, $\mathbf{y}\in\mathbb{Z}_{q}^{h}$. Then there exists a probabilistic polynomial-time (PPT) algorithm $\mathsf{GenSamplePre}(A_{S},A_{R},T_{R},\mathbf{y},V)$ that outputs a vector $\mathbf{e}\in\mathbb{Z}_{q}^{kn}$ whose distribution is statistically indistinguishable from $D_{\Lambda_{q}^{\mathbf{y}}(A_{S}),V}$.
Definition 7.
( Smoothing parameter[40]) For any $n$-dimensional lattice $\Lambda$ and any given $\epsilon>0$, the smoothing parameter of the lattice is defined as
$\eta_{\epsilon}(\Lambda)=\min\left\{s>0\mid\rho_{\frac{1}{s}}(\Lambda^{*})<1+\epsilon\right\}.$
The smoothing parameter $\eta_{\epsilon}(\Lambda_{q}^{\perp}(A))$ is at most $\omega(\sqrt{\log n})$ for some negligible $\epsilon$, and this bound holds for all but an exponentially small fraction of matrices $A\in\mathbb{Z}_{q}^{h\times n}$ [38]. The detailed description is as follows:
Lemma 4.
([38], Lemma 5.3) Let $q\geq 3$, $h$ and $n$ be positive integers satisfying $n\geq 2h\lg q$. Then there exists a negligible function $\epsilon(n)$ such that $\eta_{\epsilon}(\Lambda_{q}^{\bot}(A))<\omega(\sqrt{\log n})$ for all but at most a $q^{-h}$ fraction of $A$ in the $\mathbb{Z}_{q}^{h\times n}$ .
The following lemma demonstrates that vectors drawn from a discrete Gaussian distribution are concentrated within specific bounds with high probability.
Lemma 5.
([40]) Let $\Lambda$ be an $n$-dimensional lattice, and $T$ be a basis of the lattice $\Lambda$. If $s\geq\parallel\widetilde{T}\parallel\cdot\omega(\sqrt{\log n})$, then for any $\mathbf{c}\in\mathbb{R}^{n}$, we have:
$\mathrm{Pr}\{\parallel\mathbf{x}-\mathbf{c}\parallel>s\sqrt{n}:\mathbf{x}{\leftarrow}\mathcal{D}_{\Lambda,s,\mathbf{c}}\}\leq\mathrm{negl}(n).$
2.5 Statistical Properties Related to Discrete Gaussian Distributions
Theorem 5.
([25]) Let $\bm{t}_{i}\in\mathbb{Z}^{m}$ and $\bm{x}_{i}$ are mutually independent random variables sampled from a Gaussian distribution $D_{\bm{t}_{i}+\Lambda,\sigma}$ over $\bm{t}_{i}+\Lambda$ for $i=1,2,\cdots,k$ in which $\Lambda$ is a lattice and $\sigma\in\mathbb{R}$ is a parameter. Let $\bm{c}=(c_{1},\cdots,c_{k})\in\mathbb{Z}^{k}$ and $g=\mathrm{gcd}(c_{1},\cdots,c_{k})$, $\bm{t}=\sum_{i=1}^{k}c_{i}\bm{t}_{i}$. If $\sigma>\|\bm{c}\|\eta_{\epsilon}(\Lambda)$ for some negligible number $\epsilon$, then $\bm{z}=\sum_{i=1}^{k}c_{i}\bm{x}_{i}$ statistically closes to $D_{\bm{t}+g\Lambda,\|\bm{c}\|\sigma}$.
Lemma 6.
([24]) Let $A\in\mathbb{Z}_{q}^{n\times m},s>0,\mathbf{u}\in\mathbb{Z}^{m}$, and $\Lambda=\Lambda_{q}^{\bm{u}}(A)$. If $\mathbf{x}$ is sampled from $\mathcal{D}_{\mathbb{Z}^{m},s}$ conditioned on $A\mathbf{x}\equiv\bm{u}\pmod{q}$, then the conditional distribution of $\mathbf{x}$ is $\mathcal{D}_{\Lambda,s}$.
Lemma 7.
([24]) Let $k=\text{poly}(n)$ be even, $q=\text{poly}(n)\geq(nk)^{2}$, and $V=\sqrt{2nk\log q}\log n\geq\omega(\sqrt{\log n})$. Then $\mathcal{D}_{\mathbb{Z}^{n},V}$ and $\mathcal{D}_{\mathbb{Z}^{n},\sqrt{\frac{k\pm 2}{k}}V}$ are statistically indistinguishable.
3 Definition and Security Model of LHRS
3.1 Definition of LHRS
Definition 8.
(LHRS scheme) The linearly homomorphic ring signature is composed of a set of probabilistic polynomial-time (PPT) algorithms $\mathcal{LHRS}$ =(Setup,HR-Sign,Combine,HR-Verify), which are defined as follows:
  1. $\bullet$
    Setup ($1^{n}$, pp): Given a security parameter $n$ and public parameters pp, this algorithm outputs a key pair $(pk,sk)$. Here, the public parameters pp determine the tag space $\mathcal{T}$ , the message space $\mathcal{M}$, the signature space $\Sigma$, and and the maximum number of homomorphic operations $k_{0}$.
  2. $\bullet$
[…]
  3. $\bullet$
[…]
  4. $\bullet$
[…]
Correctness requires that:
(1) For all valid tags $\tau\in\mathcal{T}$ and messages $\mathbf{m}\in\mathcal{M}$, if a ring signature $\sigma$ is generated as $\sigma\leftarrow\textsf{HR-Sign}(\textsf{pk}_{s},\textsf{sk}_{s},R,\tau,\mathbf{m})$, then the verification algorithm must satisfy
$1\leftarrow\textsf{HR-Verify}(R,\tau,\mathbf{m},\sigma).$
(2) If $\sigma_{j}\leftarrow\textsf{HR-Sign}(\textsf{pk}_{s},\textsf{sk}_{s},R,\tau,\mathbf{m}_{j})$, then
$1\leftarrow\textsf{HR-Verify}(R,\tau,\sum_{j=1}^{\ell}c_{j}\mathbf{m}_{j},\textsf{Combine}(R,\tau,\{(c_{j},\sigma_{j})\}_{j=1}^{\ell})).$
3.2 Security Model of LHRS
In ring signature schemes [13, 15], security typically encompasses two fundamental requirements: anonymity and unforgeability. Similarly, linearly homomorphic ring signature schemes must also satisfy these two properties. Specifically, they should achieve anonymity under full key exposure and unforgeability against insider corruption.
Definition 9.
(Anonymity under Full Key Exposure ) The security of the LHRS scheme against anonymity under full key exposure is defined through the following game played between a challenger $\mathcal{B}_{1}$ and a polynomial-time adversary $\mathcal{A}_{1}$:
  1. $\bullet$
    Setup: The challenger $\mathcal{B}_{1}$ runs the $\mathsf{Setup}(1^{n},\mathsf{pp})$ algorithm $\ell$ times to generate key pairs $(\textsf{pk}_{1},\textsf{sk}_{1}),\dots,(\textsf{pk}_{\ell},\textsf{sk}_{\ell})$, where $\ell$ is a game parameter, and sends the public key set $\{\textsf{pk}_{s}\}_{s=1}^{\ell}$ to the adversary $\mathcal{A}_{1}$.
  2. $\bullet$
    Queries: After receiving $\{(\textsf{pk}_{s},\textsf{sk}_{s})\}_{s=1}^{\ell}$, the adversary $\mathcal{A}_{1}$ can make the following two types of queries:
    -Corruption query: The adversary $\mathcal{A}_{1}$ selects a user $s\in[\ell]=\{1,2,$$\ell\}$ and sends $s$ to the challenger $\mathcal{B}_{1}$, who responds by returning $\textsf{sk}_{s}$ to $\mathcal{A}_{1}$.
    -Sign query: After selecting a user $s$, adversary $\mathcal{A}_{2}$ adaptively selects a series of message subspaces $V_{si}=\text{span}\{\mathbf{m}_{si}^{(1)},...,\mathbf{m}_{si}^{(k_{0})}\}$ and queries them to the challenger $\mathcal{B}_{2}$, for $i=1,\dots,q_{s}$, where $q_{s}$ is the maximum number of allowed queries for user $s$. For each queried subspace $V_{si}$, the challenger $\mathcal{B}_{2}$ generates a uniformly random tag $\tau_{si}\xleftarrow{\mathdollar}\mathcal{T}$. Then, $\mathcal{B}_{2}$ computes a signature $\sigma_{si}^{(j)}$ for each message $\mathbf{m}_{si}^{(j)}$ within the subspace, in relation to the tag $\tau_{si}$. Finally, $\mathcal{B}_{2}$ responds with the tuples $(s,R,\tau_{si},\mathbf{m}_{si}^{(j)},\sigma_{si}^{(j)})$ for all $i=1,...,q_{s}$ and $j=1,\dots,k_{0}$.
  3. $\bullet$
    Challenge: Finally, $\mathcal{A}_{1}$ outputs a challenge tuple $(s_{0},s_{1},R^{*},\tau^{*},m^{*})$ to $\mathcal{B}_{1}$, where $s_{0}$ and $s_{1}$ are indices satisfying $\textsf{pk}_{s_{0}}\in R^{*}$ and $\textsf{pk}_{s_{1}}\in R^{*}$, and neither $s_{0}$ nor $s_{1}$ has been corrupted in any prior corruption query. $\mathcal{B}_{1}$ is then required to generate a signature on $(\tau^{*},m^{*})$ under ring $R^{*}$. To do so, $\mathcal{B}_{1}$ selects a random bit $b\xleftarrow{\mathdollar}\{0,1\}$, computes $\sigma_{b}^{*}\leftarrow\textsf{HR-Sign}(\textsf{pk}_{s_{b}},\textsf{sk}_{s_{b}},R^{*},\tau^{*},m^{*})$, and returns $\sigma_{b}^{*}$ to $\mathcal{A}_{1}$. The adversary outputs a guess $b^{\prime}\in\{0,1\}$ and wins the game if $b^{\prime}=b$.
The scheme is said to satisfy anonymity if for every probabilistic polynomial-time adversary $\mathcal{A}_{1}$, the advantage $\mathbf{Adv}_{LHRS}^{\mathrm{ANON}}(\mathcal{A}_{1})=\left|\Pr[b^{\prime}=b]-\frac{1}{2}\right|$ is negligible in the security parameter.
Definition 10.
(Unforgeability against Insider Corruption) The security of the LHRS scheme against unforgeability under insider corruption is defined through the following game played between a challenger $\mathcal{B}_{2}$ and a polynomial-time adversary $\mathcal{A}_{2}$:
  1. $\bullet$
    Setup: The challenger $\mathcal{B}_{2}$ runs the $\mathsf{Setup}(1^{n})$ algorithm $\ell$ times to generate key pairs $(\textsf{pk}_{1},\textsf{sk}_{1}),\dots,(\textsf{pk}_{\ell},\textsf{sk}_{\ell})$, where $\ell$ is a game parameter. Let $L=\{\textsf{pk}_{s}\}_{s=1}^{\ell}$ and sends the public key set $L$ to the adversary $\mathcal{A}_{2}$
  2. $\bullet$
    Queries: After receiving $L=\{\textsf{pk}_{s}\}_{s=1}^{\ell}$, the adversary $\mathcal{A}_{2}$ can make the following two types of queries:
    -Corruption query: The adversary $\mathcal{A}_{2}$ selects a user $s\in[\ell]=\{1,2,$$\ell\}$ and sends $s$ to the challenger $\mathcal{B}_{2}$, who responds by returning $\textsf{sk}_{s}$ to $\mathcal{A}_{2}$.
    -Signing query: After selecting a user $s$, adversary $\mathcal{A}_{2}$ adaptively selects a series of message subspaces $V_{si}=\text{span}\{\mathbf{m}_{si}^{(1)},...,\mathbf{m}_{si}^{(k_{0})}\}$ and queries them to the challenger $\mathcal{B}_{2}$, for $i=1,\dots,q_{s}$, where $q_{s}$ is the maximum number of allowed queries for user $s$. For each queried subspace $V_{si}$, the challenger $\mathcal{B}_{2}$ generates a uniformly random tag $\tau_{si}\xleftarrow{\mathdollar}\mathcal{T}$. Then, $\mathcal{B}_{2}$ computes a signature $\sigma_{si}^{(j)}$ for each message $\mathbf{m}_{si}^{(j)}$ within the subspace, in relation to the tag $\tau_{si}$. Finally, $\mathcal{B}_{2}$ responds with the tuples $(s,R,\tau_{si},\mathbf{m}_{si}^{(j)},\sigma_{si}^{(j)})$ for all $i=1,...,q_{s}$ and $j=1,\dots,k_{0}$.
  3. $\bullet$
    Output: Finally, $\mathcal{A}_{2}$ outputs a tuple $(R^{*},\tau^{*},\mathbf{m}^{*},\sigma^{*})$. The adversary $\mathcal{A}_{2}$ wins the game if and only if all the following conditions are satisfied:
    1. 1.
      $1\leftarrow\textsf{HR-Verify}(R^{*},\tau^{*},\mathbf{m}^{*},\sigma^{*});$
    2. 2.
      $R^{*}\subset L\setminus C$, where $\mathcal{C}$ denotes the set of all corrupted users (i.e., users whose private keys have been obtained by the adversary);
    3. 3.
      No valid signature for the tuple $(R^{*},\tau^{*},\mathbf{m}^{*})$ was generated during the signing queries in the game;
    4. 4.
      For any pair $(s,i)$, either $\tau^{*}\neq\tau_{si}$ (Type-1 forgery) where $s\in[\ell]$ and $i\in[q_{s}]$, or there exists a pair $(s,i)$ such that $\tau^{*}=\tau_{si}$ (Type-2 forgery) but $\mathbf{m}^{*}\notin V_{si}$.
The adversary $\mathcal{A}_{2}$’s advantage in the above game is defined as: $\text{Adv}_{\text{LHRS}}^{\text{UNF}}(\mathcal{A}_{2})=\Pr[\mathcal{A}_{2}\text{ wins}].$ A LHRS scheme is said to be unforgeable if for every probabilistic polynomial-time adversary $\mathcal{A}_{2}$, the advantage $\text{Adv}_{\text{LHRS}}^{\text{UNF}}(\mathcal{A}_{2})$ is negligible.
4 The Proposed Scheme
In this section, we present a lattice-based linearly homomorphic ring signature scheme and prove that the scheme satisfies correctness with overwhelming probability.
4.1 Basic Construction
The algorithms TrapGen, Decompose, and GenSamplePre used in this section are all from Section 2.
Our LHRS scheme $\mathcal{LHRS}$={Setup,HR-Sign,Combine, HR-Verify} works as follows:
  1. $\bullet$
    Setup ($1^{n}$, pp): Let $q=\text{poly}(n)$, $k_{0}=\text{poly}(n)$ (bounding maximum homomorphic operations), and $k=\text{poly}(n)$ be such that $k$ is even and $k_{0}<k$, with $q\geq(nkk_{0})^{3}$. Define $h=\left\lfloor\frac{n}{6\log q}\right\rfloor$ and $V=\sqrt{2nk\log q}\cdot\log n$. Sample $\bm{\alpha}_{1},\dots,\bm{\alpha}_{k}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{h}$, and set the public parameters as $\textsf{pp}=\{q,k_{0},k,h,V,\bm{\alpha}_{1},\dots,\bm{\alpha}_{k}\}$. The tag space is $\mathcal{T}=\{0,1\}^{k}$, the message space $\mathcal{M}=\mathbb{F}_{2}^{k}$, and the signature space $\Sigma=\mathbb{Z}^{\ell n}$ where $\ell$ denotes the number of users in ring $R$.
    The algorithm takes as input a security parameter $n$ and the public parameters pp, and generates the key pair for user $s$ as follows:
    1. (1)
      Compute $(A_{s},T_{s})\leftarrow\textsf{TrapGen}(q,h,n)$
    2. (2)
      Set the public key as $\textsf{pk}_{s}=A_{s}$ and the secret key as $\textsf{sk}_{s}=T_{s}$, then output the key pair $(\textsf{pk}_{s},\textsf{sk}_{s})=(A_{s},T_{s})$ for user $s$.
  2. $\bullet$
[…]
    1. (1)
      Let $A_{R_{\ell}}=[A_{i_{1}}|A_{i_{2}}|\cdot\cdot\cdot|A_{i_{\ell}}]$. We define an index label vector $L_{R_{\ell}}=(i_{1},i_{2},...i_{l})$ that specifies the correspondence between matrix $A_{R_{l}}$ and the cyclic member sequence $\{1,2,\ldots,\ell\}$, representing the permutation order of ring members in $A_{R_{\ell}}$.
    2. (2)
      Compute $(\mathbf{u},\mathbf{v})\leftarrow\textsf{Decompose}(\mathbf{m})$,where $\mathbf{m}=\mathbf{u}+\mathbf{v}$, $\mathbf{u}=(u_{1},...,u_{k})$ $\in\mathcal{X}$, $\mathbf{v}=(v_{1},...,v_{k})\in\mathcal{Y}$ or $\mathcal{Z}$.
    3. (3)
      If $\mathbf{u}$ and $\mathbf{v}$ contain zero vectors (note: by Theorem 1, at most one of $\mathbf{u}$ or $\mathbf{v}$ is zero), set $\mathbf{t}=\sum_{j=1}^{k}(-1)^{\tau_{j}}m_{j}\bm{\alpha}_{j}$ and compute
      $\mathbf{e}\leftarrow\mathsf{GenSamplePre}(A_{R_{\ell}},A_{s},T_{s},\mathbf{t},V);$
      otherwise, compute $\mathbf{t}(\mathbf{u})=\sum_{j=1}^{k}(-1)^{\tau_{j}}u_{j}\bm{\alpha}_{j}$ and $\mathbf{t}(\mathbf{v})=\sum_{j=1}^{k}(-1)^{\tau_{j}}v_{j}\bm{\alpha}_{j}$, then generate
      $\mathbf{e}(\mathbf{u})\leftarrow\mathsf{GenSamplePre}(A_{R_{\ell}},A_{s},T_{s},\mathbf{t}(\mathbf{u}),V),$
      $\mathbf{e}(\mathbf{v})\leftarrow\mathsf{GenSamplePre}(A_{R_{\ell}},A_{s},T_{s},\mathbf{t}(\mathbf{v}),V),$
      and define $\mathbf{e}=\mathbf{e}(\mathbf{u})+\mathbf{e}(\mathbf{v})$.
    4. (4)
      The ring signature output is denoted as $\sigma=(\mathbf{e},L_{R_{\ell}})$
  3. $\bullet$
[…]
  4. $\bullet$
[…]
    1. (1)
      Compute $A_{R_{\ell}}=[A_{i_{1}}|A_{i_{2}}|\cdot\cdot\cdot|A_{i_{\ell}}]$;
    2. (2)
      Compute $\mathbf{t}=\sum_{j=1}^{k}(-1)^{\tau_{j}}m_{j}\bm{\alpha}_{j}$
    3. (3)
      If the following three conditions are all satisfied, output 1; otherwise, output 0.
      1. 1)
        $\parallel\mathbf{e}\parallel\leq 2k_{0}V\sqrt{2k\ell n}$;
      2. 2)
        $A_{R_{\ell}}\cdot\mathbf{e}\pmod{q}=\mathbf{t}$
Remark: For the vector decomposition $(\mathbf{u},\mathbf{v})$ of a message $\mathbf{m}$, when one of the components is a zero vector, the verification of related conclusions becomes trivial; when both components are non-zero vectors, the verification process is representative of the general case. Therefore, to simplify the presentation, in proving the correctness, anonymity, and unforgeability of the scheme, we only need to analyze the case where $\mathbf{m}$ is decomposed into two non-zero vectors.
4.2 Correctness
Theorem 6.
The above $\mathcal{LHRS}$ scheme satisfies correctness with overwhelming probability.
Proof.
  1. (1)
    If $\sigma=(\mathbf{e},L_{R_{\ell}})\leftarrow\textsf{HR-Sign}(\textsf{pk}_{s},\textsf{sk}_{s},R_{\ell},\tau,\mathbf{m})$. Next, we show that the verification conditions will hold with overwhelming probability.
    By Theorems 3, we have $\parallel\widetilde{T}_{s}\parallel\leq\mathcal{O}(\sqrt{h\log q})$ with overwhelming probability. Given $V=\sqrt{2nk\log q}\cdot\log n$, we derive
    $\frac{V}{\parallel\widetilde{T}_{s}\parallel}\geq\sqrt{\frac{2kn}{h}}\cdot\log n\geq\sqrt{\log n},$
    implying $V\geq\parallel\widetilde{T}_{s}\parallel\cdot\omega(\sqrt{\log n})$. Thus, Lemma 5 ensures that, with overwhelming probability, $\parallel\mathbf{e}\parallel=\parallel\mathbf{e}(\mathbf{u})+\mathbf{e}(\mathbf{v})\parallel\leq 2V\sqrt{\ell n}<2k_{0}V\sqrt{2k\ell n}$. Moreover, Lemma 2 yields the congruence relation:
    $A_{R_{\ell}}\cdot\mathbf{e}\pmod{q}$ $=A_{R_{\ell}}\cdot(\mathbf{e}(\mathbf{u})+\mathbf{e}(\mathbf{v}))\pmod{q}$
    $=\mathbf{t}(\mathbf{u})+\mathbf{t}(\mathbf{v})$
    $=\sum_{j=1}(-1)^{\tau_{j}}u_{j}\bm{\alpha}_{j}+\sum_{j=1}(-1)^{\tau_{j}}v_{j}\bm{\alpha}_{j}$
    $=\sum_{j=1}(-1)^{\tau_{j}}(u_{j}+v_{j})\bm{\alpha}_{j}$
    $=\sum_{j=1}(-1)^{\tau_{j}}m_{j}\bm{\alpha}_{j}=\mathbf{t}.$
    Thus, for a single message signature, the scheme satisfies correctness.
  2. (2)
    Consider a collection of tuples $\{(c_{j},\sigma_{j})\}_{j=1}^{p}$, where each signature
    $\sigma_{j}=(\mathbf{e}_{j},L_{R_{\ell}})\leftarrow\textsf{HR-Sign}(\textsf{pk}_{s},\textsf{sk}_{s},R_{\ell},\tau,\mathbf{m}_{j}).$
    We demonstrate that the verification holds:
    $\textsf{HR-Verify}(R_{\ell},\tau,\sum_{j=1}^{p}c_{j}\mathbf{m}_{j},\textsf{Combine}(R_{\ell},\tau,\{(c_{j},\sigma_{j})\}_{j=1}^{p}))\rightarrow 1.$
    First, by the algorithmic definition of Combine, we obtain the following output:
    $\sigma=(\sum_{j=1}^{p}\mathbf{e}_{j},L_{R_{\ell}})\leftarrow\textsf{Combine}(R_{\ell},\tau,\{(c_{j},\sigma_{j})\}_{j=1}^{p}).$
    According to Lemma 5, we have
    $\Big\|\sum_{j=1}^{p}c_{j}\mathbf{e}_{j}\Big\|\leq 2pV\sqrt{\ell n}\leq 2k_{0}V\sqrt{\ell n}\leq 2k_{0}V\sqrt{2k\ell n}.$
    Let $\mathbf{m}=(m_{1},\dots,m_{k})=\sum_{j=1}^{p}c_{j}\mathbf{m}_{j}$, where $\mathbf{m}_{j}=(m_{j1},\dots,m_{jk})$. Then
    $m_{i}=\sum_{j=1}^{p}c_{j}m_{ji},\quad\text{for }1\leq i\leq k.$
    Since
    $A_{R_{\ell}}\cdot c_{j}\mathbf{e}_{j}(\mathrm{mod}q)=c_{j}A_{R_{\ell}}\cdot\mathbf{e}_{j}(\mathrm{mod}q)=c_{j}\mathbf{t}_{j},\quad\text{where }\mathbf{t}_{j}=\sum_{i=1}^{k}(-1)^{\tau_{i}}m_{ji}\bm{\alpha}_{i},$
    we obtain
    $A_{R_{\ell}}\cdot\left(\sum_{j=1}^{p}c_{j}\mathbf{e}_{j}\right)(\mathrm{mod}q)$ $=\sum_{j=1}^{p}c_{j}\mathbf{t}_{j}$
    $=\sum_{j=1}^{p}c_{j}\left(\sum_{i=1}^{k}(-1)^{\tau_{i}}m_{ji}\bm{\alpha}_{i}\right)$
    $=\sum_{i=1}^{k}(-1)^{\tau_{i}}\left(\sum_{j=1}^{p}c_{j}m_{ji}\right)\bm{\alpha}_{i}$
    $=\sum_{i=1}^{k}(-1)^{\tau_{i}}m_{i}\bm{\alpha}_{i}(\mathrm{mod}q)=\mathbf{t}.$
    Consequently,
    $\textsf{HR-Verify}(R_{\ell},\tau,\sum_{j=1}^{p}c_{j}\mathbf{m}_{j},\textsf{Combine}(R_{\ell},\tau,\{(c_{j},\sigma_{j})\}_{j=1}^{p}))\rightarrow 1.$
5 Anonymity and Unforgeability
5.1 Anonymity
Theorem 7.
The $\mathcal{LHRS}$ scheme constructed in Section 4 satisfies anonymity under full key exposure.
Proof.
Suppose there exists an adaptive adversary $\mathcal{A}_{1}$ that attacks our LHRS scheme according to the anonymity definition under full key exposure. Below, we construct a polynomial-time algorithm $\mathcal{B}_{1}$ to simulate the attack environment for $\mathcal{A}_{1}$.
  1. 1.
    Setup: The challenger $\mathcal{B}_{1}$ runs the $\mathsf{Setup}(1^{n},\mathsf{pp})$ algorithm $q_{E}$ times to generate key pairs $(A_{1},T_{1}),\dots,(A_{q_{E}},T_{q_{E}})$. The public parameters $\mathsf{pp}$ are identical to those defined in Section 4 of the proposed scheme. The challenger $\mathcal{B}_{1}$ stores the tuple $(s,A_{s},T_{s})$ (where $1\leq s\leq q_{E}$) in the list $L_{1}$ and provides $\{A_{1},A_{2},\ldots,A_{q_{E}}\}$ to the adversary $\mathcal{A}_{1}$.
  2. 2.
    Queries: The adversary $\mathcal{A}_{1}$ makes the following two types of queries:
    -Corruption query: The adversary $\mathcal{A}_{1}$ selects a user $s\in\{1,2,\dots,q_{E}\}$ and sends it to the challenger $\mathcal{B}_{1}$. $\mathcal{B}_{1}$ queries the list $L_{1}$ to retrieve the tuple $(s,A_{s},T_{s})$ and returns $T_{s}$ to $\mathcal{A}_{1}$.
    -Sign query: The adversary $\mathcal{A}_{1}$ selects and sends to $\mathcal{B}_{1}$ the tuple $(s,R_{\ell},V_{si})$, where the message subspace $V_{si}$ has basis vectors $\{\mathbf{m}_{si}^{(j)}\}_{j=1}^{k_{0}}$ and satisfies $\textsf{pk}_{s}\in R_{\ell}$, with $1\leq\ell\leq q_{E}$, $1\leq i\leq q_{s}$, and $q_{s}$ denoting the maximum number of queries that user $s$ can make on message subspace $V_{si}$ under ring $R_{\ell}$. $\mathcal{B}_{1}$ first generates a random tag $\tau_{si}=(\tau_{si}^{(1)},...,\tau_{si}^{(k)})\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{k}$ for $V_{si}$ and produces the signature as follows:
    1. (1)
      Compute $(\mathbf{u}_{si}^{(j)},\mathbf{v}_{si}^{(j)}){\leftarrow}\textbf{Decompose}(\mathbf{m}_{si}^{(j)})$ and let
      $\mathbf{u}_{si}^{(j)}=(u_{si}^{(j1)},...,u_{si}^{(jk)}),$
      $\mathbf{v}_{si}^{(j)}=(v_{si}^{(j1)},...,v_{si}^{(jk)}).$
    2. (2)
      Compute
      $\mathbf{t}(\mathbf{u}_{si}^{(j)})=\sum_{p=1}^{k}(-1)^{\tau_{si}^{(p)}}u_{si}^{(jp)}\bm{\alpha}_{p}$
      and
      $\mathbf{t}(\mathbf{v}_{si}^{(j)})=\sum_{p=1}^{k}(-1)^{\tau_{si}^{(p)}}v_{si}^{(jp)}\bm{\alpha}_{p}.$
    3. (3)
      Compute
      $\mathbf{e}(\mathbf{u}_{si}^{(j)})\leftarrow\textsf{GenSamplePre}(A_{R_{l}},A_{s},T_{s},\mathbf{t}(\mathbf{u}_{si}^{(j)}),V)$
      and
      $\mathbf{e}(\mathbf{v}_{si}^{(j)})\leftarrow\textsf{GenSamplePre}(A_{R_{l}},A_{s},T_{s},\mathbf{t}(\mathbf{v}_{si}^{(j)}),V).$
      Let $\mathbf{e}_{si}^{(j)}=\mathbf{e}(\mathbf{u}_{si}^{(j)})+\mathbf{e}(\mathbf{v}_{si}^{(j)})$.
    The challenger $\mathcal{B}_{1}$ returns the tuple $(s,R_{\ell},\tau_{si},\mathbf{m}_{si}^{(j)},\mathbf{e}_{si}^{(j)})$ to the adversary $\mathcal{A}_{1}$, where $\textsf{pk}_{s}\in R_{\ell}$, $\ell\in[q_{E}]$, $1\leq i\leq q_{s}$, and $1\leq j\leq k_{0}$.
  3. 3.
    Challenge: Finally, the adversary $\mathcal{A}_{1}$ outputs a challenge tuple
    $(s_{0},s_{1},R_{\ell^{*}},\tau^{*},\mathbf{m}^{*}),$
    where $s_{0}$ and $s_{1}$ are indices satisfying $\textsf{pk}_{s_{0}}\in R_{\ell^{*}}$ and $\textsf{pk}_{s_{1}}\in R_{\ell^{*}}$, and $(\tau^{*},\mathbf{m}^{*})$ is the data message to be signed by the ring $R_{\ell^{*}}$. $\mathcal{B}_{1}$ randomly selects a bit $b\in\{0,1\}$, generates a signature $\mathbf{e}_{b}^{*}$ following the same procedure as the signing query, and provides $\mathbf{e}_{b}^{*}$ to $\mathcal{A}_{1}$. Finally, $\mathcal{A}_{1}$ outputs a bit $b^{\prime}$.
Since in the environment simulated by $\mathcal{B}_{1}$ all responses are generated according to the real algorithm, the behavior of $\mathcal{B}_{1}$ is statistically close to that in the real anonymity security experiment from the adversary $A_{1}$’s perspective. Furthermore, note that $V\geq\omega(\sqrt{\log q_{E}n})\geq\omega(\sqrt{\log\ell^{*}n})\geq\eta_{\epsilon}(\Lambda_{q}^{\bot}(A_{R_{\ell^{*}}}))$. Regardless of the value of $b$, Theorem 5 shows that the distribution of $\mathbf{e}_{b}^{*}=\mathbf{e}(\mathbf{u}_{b}^{*})+\mathbf{e}(\mathbf{v}_{b}^{*})$ is statistically close to $\mathcal{D}_{\Lambda_{q}^{\mathbf{t}}(A_{R_{\ell^{*}}}),\sqrt{2}V}$, where $\mathbf{t}=\sum_{j=1}^{k}(-1)^{\tau_{j}^{*}}m_{j}^{*}\bm{\alpha}_{j}$. Therefore, the two are computationally indistinguishable.
It follows that no polynomial-time adversary $\mathcal{A}_{1}$ can distinguish between $\mathbf{e}_{0}^{*}$ and $\mathbf{e}_{1}^{*}$ with non-negligible probability, i.e.,
$\mathbf{Adv}_{LHRS}^{\mathrm{ANON}}(\mathcal{A}_{1})=\left|\Pr[b^{\prime}=b]-\frac{1}{2}\right|=\mathrm{negl}(n).$
Thus, the scheme satisfies anonymity. The proof is complete. ∎
5.2 Unforgeability
Next, we demonstrate that in schemes with smaller ring sizes (for example, less than or equal to $\mathcal{O}(\log n)$), the aforementioned $\mathcal{LHRS}$ satisfies unforgeability.
Theorem 8.
Assuming the $\mathsf{SIS}_{q,h,\ell n,\beta}$ problem is hard, where $\ell$ denotes the size of the challenge ring and $\beta=4k_{0}V\sqrt{2k\ell n}$, then the $\mathcal{LHRS}$ scheme constructed in Section 4 achieves unforgeability against against insider corruption.
More specifically, for a polynomial-time adversary $\mathcal{A}_{2}$ who is allowed to access rings of maximum size $q_{E}(\leq\mathcal{O}(\log n))$ and can make at most $q_{s}$ signing queries per user $s$ in any ring (relative to the message subspace), we can construct a polynomial-time algorithm $\mathcal{B}_{2}$ that solves the $\mathsf{SIS}_{q,h,\ell n,\beta}$ problem with the following advantage:
$\mathbf{Adv}_{q,h,\ell n,2V}^{\mathrm{SIS}}(\mathcal{B}_{2})\geq\frac{\mathbf{Adv}_{\mathrm{LHRS}}^{\mathrm{UNF}}(\mathcal{A}_{2})}{2q_{E}\binom{q_{E}}{q_{E}/2}}-\mathrm{negl}(n).$
Proof.
Suppose the challenger $\mathcal{B}_{2}$ receives a challenge instance $A_{R_{\mathbf{w}}}\in\mathbb{Z}_{q}^{h\times\ell n}$ of the $\mathsf{SIS}_{q,h,\ell n,\beta}$ problem, with the goal of finding a nonzero short vector $\mathbf{e}_{0}$ that satisfies $A_{R_{\mathbf{w}}}\mathbf{e}_{0}\equiv\mathbf{0}\pmod{q}$ and $\|\mathbf{e}_{0}\|\leq\beta$.
$\mathcal{B}_{2}$ knows that the adversary $\mathcal{A}_{2}$ can perform at most $q_{E}$ key generation operations (i.e., the system has at most $q_{E}$ users), and that each user $s$ can make at most $q_{s}$ signature queries (relative to the number of message subspaces). $\mathcal{B}_{2}$ attempts to guess which ring $R^{*}$ $\mathcal{A}_{2}$ will ultimately choose for forgery and its size $\ell$. It randomly selects an $\ell\in[q_{E}]$ (guessing the ring size) and an index vector $\mathbf{w}=(w_{1},...,w_{\ell})\in[q_{E}]^{\ell}$ (guessing the ring member indices), setting the target ring as $R_{\mathbf{w}}=\{A_{w_{1}},...,A_{w_{\ell}}\}$.
  1. $\bullet$
    Setup: In the system, the parameters $q$, $k$, $h$, and the size of $V$ follow the same configuration as specified in Section 4. Sample $k$ vectors: $\mathbf{x}_{1},\dots,\mathbf{x}_{k}\xleftarrow{\mathdollar}\textsf{SampleDom}(1^{\ell n},\frac{1}{\sqrt{k/2}}V)$. Let $\bm{\alpha}_{1}=A_{R_{\mathbf{w}}}\mathbf{x}_{1}$, $\dots$, $\bm{\alpha}_{k}=A_{R_{\mathbf{w}}}\mathbf{x}_{k}$. Observe that $\frac{1}{\sqrt{k/2}}V\geq\omega\big(\sqrt{\log(q_{E}n)}\big)\geq\omega\big(\sqrt{\log(\ell n)}\big)$. According to Lemma 3, the vectors $\bm{\alpha}_{1},\dots,\bm{\alpha}_{k}$ are statistically close to the uniform distribution over $\mathbb{Z}_{q}^{h}$. $\mathcal{B}_{2}$ then sends the public parameters $\textsf{pp}=\{q,k_{0},k,h,V,\bm{\alpha}_{1},\dots,\bm{\alpha}_{k}\}$ to $\mathcal{A}_{2}$. Finally, $\mathcal{B}_{2}$ generates public-private key pairs for all $q_{E}$ users in the following manner:
    1. Case 1: $s\notin\mathbf{w}$. The $\mathcal{B}_{2}$ invokes the algorithm $\textsf{TrapGen}(q,h,n)$ to generate the key pair $(A_{s},T_{s})$ for user $s$, where $A_{s}$ and $T_{s}$ are the public and private keys, respectively. The tuple $(s,A_{s},T_{s})$ is stored in the list $L_{2}$, initially empty.
    2. Case 2: $s\in\mathbf{w}$. Assume that $s=w_{s^{*}}$, meaning user $s$ is the $s^{*}$-th member of the target ring $R_{\mathbf{w}}$. Representing the matrix $A_{R_{\mathbf{w}}}$ in column-block form as $A_{R_{\mathbf{w}}}=[A_{w_{1}}|A_{w_{2}}|\cdot\cdot\cdot|A_{w_{\ell}}]$, $\mathcal{B}_{2}$ assigns the $s^{*}$-th block $A_{w_{s^{*}}}$ of the SIS instance matrix $A_{R_{\mathbf{w}}}$ to user $s$, and sets their public key as $A_{s}=A_{w_{s^{*}}}$. Finally, $\mathcal{B}_{2}$ stores $(s,A_{s})$ in list $L_{2}$.
    The challenger $\mathcal{B}_{2}$ sends all users’ public keys $L=\{A_{1},\ldots,A_{q_{E}}\}$ to the adversary $\mathcal{A}_{2}$.
  2. $\bullet$
    Query: $\mathcal{A}_{2}$ makes the following two types of queries:
    -Corruption query: When $\mathcal{A}_{2}$ queries the private key of user $s$, $\mathcal{B}_{2}$ responds as follows:
    1. Case 1: $s\notin\mathbf{w}$. $\mathcal{B}_{2}$ searches the list $L_{2}$ for the entry $(s,A_{s},T_{s})$ and sends the corresponding $T_{s}$ to $\mathcal{A}_{2}$.
    2. Case 2: $s\in\mathbf{w}$. $\mathcal{B}_{2}$ aborts.
    -Signing query : $\mathcal{A}_{2}$ selects the tuple $(s,R_{\mathbf{w}^{*}},V_{si},)$ and sends it to $\mathcal{B}_{2}$, where the basis vectors of the message subspace $V_{si}$ are $\{\mathbf{m}_{si}^{(j)}\}_{j=1}^{k}$, $1\leq i\leq q_{s}$. $\mathcal{B}_{2}$ randomly generates a tag $\tau_{si}=(\tau_{si}^{(1)},...,\tau_{si}^{(p)})\in\{0,1\}^{k}$ for each message subspace $V_{si}$. $\mathcal{B}_{2}$ constructs the signature according to different cases:
    1. Case 1: $R_{\mathbf{w}^{*}}=R_{\mathbf{w}}$.
      1. (1)
        Compute $(\mathbf{u}_{si}^{(j)},\mathbf{v}_{si}^{(j)}){\leftarrow}\textbf{Decompose}(\mathbf{m}_{si}^{(j)})$ and let
        $\mathbf{u}_{si}^{(j)}=(u_{si}^{(j1)},...,u_{si}^{(jk)}),$
        $\mathbf{v}_{si}^{(j)}=(v_{si}^{(j1)},...,v_{si}^{(jk)}).$
      2. (2)
        Compute
        $\mathbf{e}(\mathbf{u}_{si}^{(j)})=\sum_{p=1}^{k}(-1)^{\tau_{si}^{(p)}}u_{si}^{(jp)}\mathbf{x}_{p}$
        and
        $\mathbf{e}(\mathbf{u}_{si}^{(j)})=\sum_{p=1}^{k}(-1)^{\tau_{si}^{(p)}}u_{si}^{(jp)}\mathbf{x}_{p}.$
        Let $\mathbf{e}_{si}^{(j)}=\mathbf{e}(\mathbf{u}_{si}^{(j)})+\mathbf{e}(\mathbf{v}_{si}^{(j)})$
      3. (3)
        $\mathcal{B}_{2}$ returns the tuple $(s,R_{\mathbf{w}^{*}},\tau_{si},\mathbf{m}_{si}^{(j)},\mathbf{e}_{si}^{(j)})$ to $\mathcal{A}_{2}$.
    2. Case 2: $R_{\mathbf{w}^{*}}\neq R_{\mathbf{w}}$, and $s\notin\mathbf{w}$.
      1. (1)
        $\mathcal{B}_{2}$ retrieves the tuple $(s,A_{s},T_{s})$ from the list $L_{2}$ to obtain the private key $T_{s}$ of user $s$.
      2. (2)
        Compute $(\mathbf{u}_{si}^{(j)},\mathbf{v}_{si}^{(j)}){\leftarrow}\textbf{Decompose}(\mathbf{m}_{si}^{(j)})$ and let
        $\mathbf{u}_{si}^{(j)}=(u_{si}^{(j1)},...,u_{si}^{(jk)}),$
        $\mathbf{v}_{si}^{(j)}=(v_{si}^{(j1)},...,v_{si}^{(jk)}).$
      3. (3)
        Compute
        $\mathbf{t}(\mathbf{u}_{si}^{(j)})=\sum_{p=1}^{k}(-1)^{\tau_{si}^{(p)}}u_{si}^{(jp)}\bm{\alpha}_{p}$
        and
        $\mathbf{t}(\mathbf{v}_{si}^{(j)})=\sum_{p=1}^{k}(-1)^{\tau_{si}^{(p)}}v_{si}^{(jp)}\bm{\alpha}_{p}.$
      4. (4)
        Compute
        $\mathbf{e}(\mathbf{u}_{si}^{(j)})\leftarrow\textsf{GenSamplePre}(A_{R_{\mathbf{w}^{*}}},A_{s},T_{s},\mathbf{t}(\mathbf{u}_{si}^{(j)}),V)$
        and
        $\mathbf{e}(\mathbf{v}_{si}^{(j)})\leftarrow\textsf{GenSamplePre}(A_{R_{\mathbf{w}^{*}}},A_{s},T_{s},\mathbf{t}(\mathbf{v}_{si}^{(j)}),V).$
        Let $\mathbf{e}_{si}^{(j)}=\mathbf{e}(\mathbf{u}_{si}^{(j)})+\mathbf{e}(\mathbf{v}_{si}^{(j)})$.
      5. (5)
        $\mathcal{B}_{2}$ returns the tuple $(s,R_{\mathbf{w}^{*}},\tau_{si},\mathbf{m}_{si}^{(j)},\mathbf{e}_{si}^{(j)})$ to $\mathcal{A}_{2}$.
    3. Case 3: $R_{\mathbf{w}^{*}}\neq R_{\mathbf{w}}$, and $s\in\mathbf{w}$.
      1. (1)
        $\mathcal{B}_{2}$ searches for a member $s^{*}$ in the ring $R_{\mathbf{w}^{*}}$ that satisfies $s^{*}\notin\mathbf{w}$, then retrieves the tuple $(s^{*},A_{s^{*}},T_{s^{*}})$ from the list $L_{2}$ to obtain the private key $T_{s^{*}}$ of user $s^{*}$.
      2. (2)
        The remaining steps follow the signature generation procedure of Case 2 to produce the signature $\mathbf{e}_{si}^{(j)}$. Note that any member of the ring $R_{\mathbf{w}^{*}}$ can generate valid signatures on behalf of this ring.
      3. (3)
        $\mathcal{B}_{2}$ returns the tuple $(s,R_{\mathbf{w}^{*}},\tau_{si},\mathbf{m}_{si}^{(j)},\mathbf{e}_{si}^{(j)})$ to $\mathcal{A}_{2}$.
We now demonstrate that the signatures provided by $\mathcal{B}_{2}$ are both valid and indistinguishable from those in the real scheme.
Since the signatures in both Case 2 and Case 3 are generated by the real algorithm, we only need to demonstrate that the signatures produced in Case 1 are both valid and computationally indistinguishable from genuine signatures.
From Theorem 5 and Lemma 7, it follows that the distributions of both $\mathbf{e}(\mathbf{u}_{si}^{(j)})$ and $\mathbf{e}(\mathbf{v}_{si}^{(j)})$ are statistically close to the distribution $\mathcal{D}_{\mathbf{Z}^{\ell^{*}n},V}$, where $\ell^{*}$ denotes the number of ring members in $R_{\mathbf{w}^{*}}$.
Furthermore, since $A_{R_{\mathbf{w}^{*}}}\cdot\mathbf{e}(\mathbf{u}_{si}^{(j)})=\mathbf{t}(\mathbf{u}_{si}^{(j)})\pmod{q}$, where $\mathbf{t}(\mathbf{u}_{si}^{(j)})=\sum_{p=1}^{k}(-1)^{\tau_{si}^{(p)}}u_{si}^{(jp)}\alpha_{p}$, it follows from Lemma 6 that $\mathbf{e}(\mathbf{u}_{si}^{(j)})$ is statistically close to the distribution $\mathcal{D}_{\Lambda_{q}^{\mathbf{t}(\mathbf{u}_{si}^{(j)})},V}(A_{R_{\mathbf{w}^{*}}})$. Similarly, $\mathbf{e}(\mathbf{v}_{si}^{(j)})$ is statistically close to the distribution $\mathcal{D}_{\Lambda_{q}^{\mathbf{t}(\mathbf{v}_{si}^{(j)})},V}(A_{R_{\mathbf{w}^{*}}})$.
According to Lemma 5, we have
$\left\|\mathbf{e}(\mathbf{u}_{si}^{(j)})\right\|=\left\|\sum_{p=1}^{k}(-1)^{\tau_{si}^{(p)}}u_{si}^{(jp)}\mathbf{x}_{p}\right\|\leq k\max_{1\leq p\leq k}\left\|\mathbf{x}_{p}\right\|\leq k\frac{V}{\sqrt{k/2}}\sqrt{\ell^{*}n}=V\sqrt{2k\ell^{*}n}.$
Similarly,
$\left\|\mathbf{e}(\mathbf{v}_{si}^{(j)})\right\|\leq V\sqrt{2k\ell^{*}n}.$
Consequently,
$\left\|\mathbf{e}_{si}^{(j)}\right\|=\left\|\mathbf{e}(\mathbf{u}_{si}^{(j)})+\mathbf{e}(\mathbf{v}_{si}^{(j)})\right\|\leq 2V\sqrt{2k\ell^{*}n}\leq 2k_{0}V\sqrt{2k\ell^{*}n}.$
Meanwhile, the adversary $\mathcal{A}_{2}$ can verify the following equation holds:
$A_{R_{\mathbf{w}^{*}}}\cdot\mathbf{e}_{si}^{(j)}$ $=A_{R_{\mathbf{w}^{*}}}\cdot\left(\mathbf{e}(\mathbf{u}_{si}^{(j)})+\mathbf{e}(\mathbf{v}_{si}^{(j)})\right)$
$=A_{R_{\mathbf{w}^{*}}}\left(\sum_{p=1}^{k}(-1)^{\tau_{si}^{(p)}}u_{si}^{(jp)}\mathbf{x}_{p}+\sum_{p=1}^{k}(-1)^{\tau_{si}^{(p)}}v_{si}^{(jp)}\mathbf{x}_{p}\right)$
$=\sum_{p=1}^{k}(-1)^{\tau_{si}^{(p)}}u_{si}^{(jp)}A_{R_{\mathbf{w}^{*}}}\cdot\mathbf{x}_{p}+\sum_{p=1}^{k}(-1)^{\tau_{si}^{(p)}}v_{si}^{(jp)}A_{R_{\mathbf{w}^{*}}}\cdot\mathbf{x}_{p}$
$=\sum_{p=1}^{k}(-1)^{\tau_{si}^{(p)}}u_{si}^{(jp)}\bm{\alpha}_{p}+\sum_{p=1}^{k}(-1)^{\tau_{si}^{(p)}}v_{si}^{(jp)}\bm{\alpha}_{p}$
$=\sum_{p=1}^{k}(-1)^{\tau_{si}^{(p)}}\left(u_{si}^{(jp)}+v_{si}^{(jp)}\right)\bm{\alpha}_{p}$
$=\sum_{p=1}^{k}(-1)^{\tau_{si}^{(p)}}m_{si}^{(jp)}\bm{\alpha}_{p}$
$=\mathbf{t}_{si}^{(j)}\pmod{q}$
Therefore, the signatures simulated by $\mathcal{B}_{2}$ are both valid and computationally indistinguishable from those in the real scheme.
After completing the queries, $\mathcal{A}_{2}$ outputs a valid forgery $(R^{*},\tau^{*},\mathbf{m}^{*},\sigma^{*})$, where $\sigma^{*}=(\mathbf{e}^{*},L_{R^{*}})$ and $L_{R^{*}}$ denotes the index of the ring $R^{*}$.
If $R^{*}\neq R_{\mathbf{w}}$, then $\mathcal{B}_{2}$ aborts the simulation. Otherwise, for either Type 1 or Type 2 forgery, the verification condition implies:
$A_{R_{\mathbf{w}}}\cdot\mathbf{e}^{*}$ $=\mathbf{t}^{*}=\sum_{i=1}^{k}(-1)^{\tau_{i}^{*}}m_{i}^{*}\bm{\alpha}_{i}$
$=A_{R_{\mathbf{w}}}\sum_{i=1}^{k}(-1)^{\tau_{i}^{*}}m_{i}^{*}\mathbf{x}_{i}\pmod{q}$.
Consequently,
$A_{R_{\mathbf{w}}}\cdot\left(\mathbf{e}^{*}-\mathbf{x}\right)\equiv\mathbf{0}\pmod{q}$,
where $\mathbf{x}=\sum_{i=1}^{k}(-1)^{\tau_{i}^{*}}m_{i}^{*}\mathbf{x}_{i}$.
Moreover, since
$\left\|\mathbf{x}\right\|=\left\|\sum_{i=1}^{k}(-1)^{\tau_{i}^{*}}m_{i}^{*}\mathbf{x}_{i}\right\|\leq k\cdot\frac{V}{\sqrt{k/2}}\sqrt{\ell n}=V\sqrt{2k\ell n}\leq 2k_{0}V\sqrt{2k\ell n},$
it follows that
$\left\|\mathbf{e}^{*}-\mathbf{x}\right\|\leq 4k_{0}V\sqrt{2k\ell n}.$
If $\mathbf{e}^{*}-\mathbf{x}\neq\mathbf{0}$, then $\mathcal{B}_{2}$ directly outputs $\mathbf{e}^{*}-\mathbf{x}$ as a valid solution to the $\mathsf{SIS}_{q,h,\ell n,\beta}$ problem.
Probability analysis: According to the properties of the sampling algorithm, since the preimages have a conditional min-entropy $\omega(\log n)$, thus
$\Pr\left[\mathbf{e}^{*}=\mathbf{x}\right]\leq 2^{-\omega(\log n)}\leq\text{negl}(n).$
Next, the probability that $\mathcal{B}_{2}$ does not abort during the game is at least $\frac{1}{2q_{E}\binom{q_{E}}{q_{E}/2}}$.
Therefore, $\mathcal{B}_{2}$ will obtain a non-zero short solution to the $\mathsf{SIS}_{q,h,\ell n,\beta}$ problem with the following advantage:
$\mathbf{Adv}_{q,h,\ell n,2V}^{\mathrm{SIS}}(\mathcal{B}_{2})\geq\frac{\mathbf{Adv}_{\mathrm{LHRS}}^{\mathrm{UNF}}(\mathcal{A}_{2})}{2q_{E}\binom{q_{E}}{q_{E}/2}}-\mathrm{negl}(n).$
Analysis of the Security-Loss Factor. The assumption $q_{E}\leq\mathcal{O}(\log n)$ implies that the binomial coefficient $\binom{q_{E}}{q_{E}/2}$ is at most polynomial in $n$. Specifically, using the approximation $\binom{2k}{k}\sim 4^{k}/\sqrt{\pi k}$, the entire denominator $2q_{E}\cdot\binom{q_{E}}{q_{E}/2}$ is of order $\mathcal{O}(\mathsf{poly}(n))$. Therefore, the reduction only incurs a polynomial loss in the advantage. A non-negligible success probability for the forger $\mathcal{A}_{2}$ thus translates to a non-negligible advantage for the SIS solver $\mathcal{B}_{2}$.
Worst-case to Average-case Reduction. Building upon Theorem 2 in [38], the computational complexity of the $\mathrm{SIS}_{q,n,m,\beta}$ can be shown to be polynomially equivalent to worst-case approximations of the Shortest Independent Vectors Problem (SIVP), given that the modulus satisfies $q\geq\beta\cdot\omega(\sqrt{n\log n})$. Under this condition, the approximation factor for SIVP is on the order of $\beta\cdot\tilde{O}(\sqrt{n})$. In our construction, we require that $q\geq(nkk_{0})^{3}$. Therefore, there exists a sufficiently large $q$ that meets the conditions of Theorem 2, thereby ensuring its applicability.
6 Conclusions
This work presents the first lattice-based linearly homomorphic ring signature (LHRS) scheme, combining the anonymity of ring signatures with the computational flexibility of homomorphic signatures. Our construction achieves quantum-resistant security while maintaining practical efficiency, addressing a significant gap in cryptographic primitives. The scheme’s unconditional anonymity and existential unforgeability make it suitable for privacy-preserving applications, such as confidential blockchain transactions and secure multi-party computation.
This scheme still presents several promising directions for further research and improvement: (1) First, explore the construction of novel schemes featuring short signatures (where the signature size is independent of the ring size) with reasonable computational efficiency; (2) Investigating the feasibility of designing homomorphic ring signature schemes that support polynomial or fully homomorphic computations; (3) Examining whether homomorphic ring signature schemes with tight or almost-tight security can be constructed under standard lattice assumptions.
Acknowledgments
We gratefully acknowledge the support and encouragement of this work from the Academy of Interdisciplinary Studies, Renmin University of China, and Great Bay University.
Author contributions
These authors contributed equally to this work.
Funding
This work was supported by Information Security School-Enterprise Joint Laboratory(Dongguan Institute for Advanced Study, Great Bay University, NO.H24120002).
Availability of data and materials
No data were used in the present study. Thus, there is no relevant data to share or report regarding availability.
Declarations
The authors declare that they have no known competing financial interests or personal relationships that could have influenced the work reported in this paper. All authors confirm that the research complies with ethical standards and is free from conflicts of interest.